Carlsbad Corporation's sales are expected to increase from $5 million in 2019 to $6 million in 2020, or by 20%. Its assets totaled $4 million at the end of 2019. Carlsbad is at full capacity, so its assets must grow in proportion to projected sales. At the end of 2019, current liabilities are $1 million, consisting of $250,000 of accounts payable, $500,000 of notes payable, and $250,000 of accrued liabilities. Its profit margin is forecasted to be 5%, and the forecasted retention ratio is 45%. Use the AFN equation to forecast the additional funds Carlsbad will need for the coming year. Write out your answer completely. For example, 5 million should be entered as 5,000,000. Round your answer to the nearest dollar.

$  
Answer #1

Percentage increase in sales= 20%

2020 retained earnings= net income margin* sales* retention ratio

= 5%*6000000*45%

= 135000

AFN = Increase in assets-Increase in spontaneous liabilities -Retained earnings

Increase in assets =4000000*20% =800000

Increase in spontaneous liabilities= (250000+250000)*20%=100000

AFN= 800000-100000-135000

AFN =$565000